x y 1 5.92 2 9.44 3 11.36 4 14.08 5 17 6 15.92 7 17.74...

x y
1 5.92
2 9.44
3 11.36
4 14.08
5 17
6 15.92
7 17.74
8 22.56
9 23.88
10 25.3
11 27.32
12 28.04
13 32.76
14 34.28

based on the data shown, calculate the regression line (each value to two decimal places)

y= __________x + _______
Answer #1

Sum of X = 105
Sum of Y = 285.6
Mean X = 7.5
Mean Y = 20.4
Sum of squares (SSX) = 227.5
Sum of products (SP) = 469.1

Regression Equation = ŷ = bX + a

b = SP/SSX = 469.1/227.5 = 2.06

a = MY - bMX = 20.4 - (2.06*7.5) = 4.94

ŷ = 2.06X + 4.94